OpenStudy (anonymous):

(s+r)/(sr^2 )+(9s+r)/(s^2 r)

OpenStudy (anonymous):

\[\left(\begin{matrix}s+r \\ sr ^{2}\end{matrix}\right)+\left(\begin{matrix}9s+r \\ s ^{2}r\end{matrix}\right)\]